### v1

# references.bib

Abstract.
Recent advancements in commodity server processors have enabled dynamic hardware-based quality-of-service (QoS) enforcement. These features have gathered increasing interest in research communities due to their versatility and wide range of applications. Thus, there exists a need to understand how scholars leverage hardware QoS enforcement in research, understand strengths and shortcomings, and identify gaps in current state-of-the-art research. This paper observes relevant publications, presents a novel taxonomy, discusses the approaches used, and identifies trends. Furthermore, an opportunity is recognized for QoS enforcement utilization in service-based cloud computing environments, and open challenges are presented.
Hardware QoS Enforcement, Cache Allocation, Resource Management, Cloud Computing, Congestion Control, Kubernetes
copyright: ccccs: Computer systems organization Cloud computingccs: Computer systems organization Real-time systemsccs: Software and its engineering Memory managementccs: Hardware Semiconductor memory
1. Introduction
Support for management and partitioning of shared system resources such as shared cache and memory bandwidth has recently become available in commercial off-the-shelf x86 server processors. The features, which we refer to as hardware-based quality-of-service (QoS) enforcement features, may be dynamically controlled at runtime using processor model-specific registers or an operating system-level interface. They enable systems to enforce availability guarantees and mitigate the interference caused by shared resource congestion. The benefits of intelligently tuning such features have been shown in previous work to include increased system performance (lo_heracles_2015), improved system utilization (patel_clite_2020), and reduced power consumption (nejat_cooperative_2022). However, to the best of our knowledge, there has been no previous attempt at surveying the academic and industrial landscapes to understand the applicability of these new features in different contexts.
One such context of interest is that of general-purpose cloud computing. The cloud-connected nature of day-to-day applications implies that performance issues of cloud-based web services now have a prevalent impact on the daily lives of the general population in a way that was unthinkable a few decades ago. In fact, a (ballard_industry_2023) US-based survey indicated that $48\%$ of companies use industry-tailored clouds with another $20\%$ intending to make the transition to cloud-based solutions within 12 months (ballard_industry_2023). Latency-sensitive web services are often referred to as latency critical (LC) tasks, which is one of two broad categories cloud-based workloads may be classified into. The response times of LC services often directly affect the end-user experience and are therefore monitored and kept to performance targets specified by service level objectives (SLOs). Best-effort (BE) jobs on the other hand are time non-critical, predictable workloads that may run whenever sufficient resources become available, commonly using a batch-format scheduling system. Thus, such jobs are well suited to improve server utilization by consuming resources left unused by LC services based on current loads, a technique leveraged by several previous works (lo_heracles_2015; chen_parties_2019; nikas_dicer_2019; patel_clite_2020; papadakis_improving_2017; qin_locpart_2017; qiu_slo_2021; zhao_rhythm_2020; li_rldrm_2020; zhu_dirigent_2016; chen_avalon_2019). BE tasks such as machine learning model training or big data analysis jobs are often data intensive and consequently disproportionately “noisy” users of shared system resources. Unless mitigated, this congestion may negatively affect the performance of LC services (dasari_identifying_2013) and subsequently, end user experience.
The congestion of shared resources such as last-level cache (LLC) and memory bandwidth (MB) caused by the co-location of workloads on multicore systems has been shown in previous studies to reduce system performance (dasari_identifying_2013; lofwenmark_understanding_2016; zhuravlev_addressing_2010). The importance of this becomes evident when considering the SLOs that web services often are required to meet. Not reaching performance targets due to resource congestion may end up costly as BE jobs may need to be canceled and re-queued. Alternatively, additional computing resources may need to be provisioned if permitted by the underlying infrastructure platform. Hardware-based QoS enforcement features are designed to mitigate the issues of shared resource congestion. However, hardware-based QoS enforcement is indirect in nature (see Section 2) and thus, intelligent tuning is required to achieve predictable and performant service operations in cloud environments.
Although intelligent management of shared resources has been shown to improve workload performance and energy efficiency, popular hypervisors and container orchestration engines still lack support for autonomous management of QoS enforcement on supported platforms. We thus identify an opportunity for further performance improvements in general-purpose cloud computing environments, but also question why the adoption of QoS enforcement has not yet been made widespread despite the promises of improved performance. Consequently, there exists a need to review the approaches proposed by researchers that utilize hardware-based QoS enforcement features. We review published research and map how the features provide value and the problems they help solve. Furthermore, we propose a novel taxonomy and investigate the contexts in which such features are discussed such that gaps in current state-of-the-art research and areas of feature application may be identified. Additionally, a discussion of the reasons behind the lack of autonomous QoS enforcement management solutions available in widespread workload management systems is presented.
1.1. Our Contributions
In this survey, we perform a thorough review of published research that integrates hardware-based QoS enforcement features as an integral part of their proposals. Thus, a complete scope of the currently identified applicability of features may be outlined. In summary, we make the following contributions:
  1. (1)
    We survey a search space of $3685$ articles and identify $79$ papers that fit the inclusion criteria defined for our review.
  2. (2)
    We present a novel taxonomy of the included state-of-the-art research that details approaches and techniques that leverage hardware QoS enforcement.
  3. (3)
    We discuss the ways hardware QoS enforcement features are utilized and identify reasons why they have not reached widespread adoption in autonomous cloud computing environments.
  4. (4)
    We detail gaps in current state-of-the-art research and present opportunities for future investigations.
2. Hardware-based QoS Enforcement Features and their Origins
The rapidly increasing number of cores available on a single multi-core processor has enabled larger numbers of workloads to be co-located on a single machine. While this has its advantages, it puts increasing pressure on the resources shared among the cores of a system. Previous studies (dasari_identifying_2013; zhuravlev_addressing_2010; lofwenmark_understanding_2016) have identified the main sources of performance interference in multi-core systems. dasari_identifying_2013 (dasari_identifying_2013) identified several main sources of unpredictable delays due to resource interference. They conclude that shared main memory, shared cache, and shared interconnection networks are the main sources of unpredictability. These claims are supported by lofwenmark_understanding_2016 (lofwenmark_understanding_2016). Furthermore, zhuravlev_addressing_2010 (zhuravlev_addressing_2010) found in their (zhuravlev_addressing_2010) study that shared cache, main memory bus, and memory access controller congestion were all prominent sources of interference.
To address the problem of shared cache interference, iyer_cqos_2004 introduced the notion of QoS in the context of shared cache management in (iyer_cqos_2004) (iyer_cqos_2004). They theorized an early cache partitioning approach that was evaluated using cache simulators. While this proposal was rudimentary, it was developed upon and more refined proposals of cache and memory partitioning approaches were later published (iyer_qos_2007). Simultaneously, low-overhead shared cache monitoring solutions were also theorized (zhao_cachescouts_2007).
In more recent years, chip manufacturers have started to implement cache and memory bandwidth management features in hardware. Intel refers to these features as Intel Resource Director Technology (RDT) (intel_corporation_intel_2019), while AMD calls their implementation AMD Platform Quality of Service Extensions (QoSE) (advanced_micro_devices_inc_amd64_2022). The features allow practitioners to tune hardware to the specific needs of their workloads and thus achieve improved efficiency through reduced interference.
This section discusses the causes of performance unpredictability originating in shared resource congestion. This requires an insight into the architecture of modern multi-core processors, as well as how they implement the specific resource of interest. Additionally, the inner workings of the dynamic resource control mechanisms of modern processors are discussed.
2.1. Classes of Service
The control mechanisms of hardware-based QoS enforcement features use a concept called classes of service (COS or CLOS). This concept is grounded in hardware design and dictates how resource control rules are applied to the workloads of a system. Every logical core is assigned to one, and only one, COS at any one time. Resource control rules are set at the COS level, which will be enforced upon any core in the COS. As the system starts up, all cores are assigned to COS0, the default COS (yu_24_nodate). They may later be assigned to different COSs dynamically at runtime. The total number of available COSs is dependent on the processor model and thus changes between systems.
2.2. Shared Cache Congestion
Traditional von Neumann architecture systems have processing and memory units physically separated (arikpo2007neumann). Thus, the performance of a system becomes highly dependent on the performance of the interconnect between the two physical units, which has become known as the von Neumann bottleneck (arikpo2007neumann). Over the past few decades, advances in processing technology have increasingly outgrown that of memory technology (hennessy_computer_2011) and this difference in performance growth is referred to as the memory wall (iyer_advances_2021), which has prompted the invention of many techniques to mitigate the problem.
One such technique is to use a cache. A cache is a small, high-performance memory package that is typically embedded in the CPU die. As the cache is substantially faster than the main memory, the processor may place data currently worked on in the cache and thus spend less time waiting for the data it needs from the main memory (brett_memory_2016).
The caches of modern multi-core processors are so-called multi-layer caches. Each layer of cache is larger, but slower than the last. Typically, modern caches consist of three layers. Layer 1 (L1) is the smallest and fastest layer. Each physical core has a separate L1 cache split into two parts, a data cache (L1d), and an instruction cache (L1i). The next layer, layer 2 (L2), is similarly constructed so that each physical core has its own instance. However, there is no distinction between data and instruction memory in the L2 cache. Layer 3 (L3) is shared among multiple cores, often all cores on a chip. This means that there may be inter-core interference as the workloads of different cores compete for cache space. The highest numbered level of cache on a system is commonly known as the last-level cache (LLC). On a three-level cache system, LLC refers to the L3 cache. An overview of a three-layer cache architecture is illustrated in Figure 1.
Figure 1. Illustration of a typical three-layer cache architecture.
Due to the cache’s small size, it is inevitable that memory will have to be read from the main memory eventually as working set sizes outgrow the available cache. AMD’s EPYC Genoa server processors with 3D V-Cache technology provide up to $1152$ megabytes of LLC cache while supporting up to 6 terabytes of main memory per socket (advanced_micro_devices_inc_amd_2023) and thus, LLC accounts for only $0.02\%$ of the main memory size. When memory required by the processor is not already present in the cache, a cache miss has occurred and a delay penalty is incurred while waiting for the required data (dasari_identifying_2013). In a (velten_memory_2022) study by velten_memory_2022, it was found that L3 access incurred a penalty of 39 cycles while waiting for main memory access took 221 cycles (velten_memory_2022)111 L3 and main memory access times are highly dependent on hardware (velten_memory_2022). The numbers presented were recorded on a system equipped with two AMD EPYC 7702 (Rome) processors. . Thus, accessing data not present in any level of cache results in a 5.67×5.67$\times$5.67 × increase in access time compared to accessing data present in the shared L3 cache.
Modern data centers spend an increasing proportion of computational resources on data-intensive workloads such as machine learning (ML) training, ML inference, video processing, and audio processing. The size of the data sets used by these processes means that they are particularly destructive to cache integrity, meaning more time is spent waiting for data from the main memory. Given the shared nature of the LLC, this affects all workloads of a system and is known as shared cache congestion. Such congestion may take the form of inter-task interference when a process replaces cache lines of another process on the same core, or inter-core interference when cache lines of a process running on another core are overridden (lofwenmark_understanding_2016).
2.2.1. Shared Cache Partitioning
Hardware support for cache partitioning has been implemented by chip manufacturers in recent years. Tuning the cache partitioning scheme may reduce the impact of shared cache congestion on workloads. As we will see in this survey, it has been repeatedly shown that cache partitioning may be used to great effect when optimizing the workloads of a system.
Configuration of the cache allocation scheme for a COS is done by setting a bitmask, the size of which is dependent on the specific CPU model. A CPU supporting $n$ bits of control may have a COS assigned $1/n$ or more [^2] COSs assigned more than one such share must on Intel platforms have all “1” bits in a contiguous block due to hardware limitations (intel_corporation_intel_2023). of the available cache. Allocation bitmasks may overlap between COSs. If so, parts of the cache are shared between two or more classes. An illustration of a cache partitioning configuration may be observed in Figure 2.
Figure 2. Illustration of a cache partitioning scheme for a CPU supporting $8$ bits of control ($0_{16}$ to $\textrm{FF}_{16}$). Here, COS0 has exclusive access to $2/8=1/4$ of available cache, while COS1 and COS2 share parts of their allocation.
When moving a core to a new COS, the new rules only apply to future cache allocations. Cache lines already present in the old partition may continue to be used until evicted (yu_24_nodate).
Some hardware additionally supports the code and data prioritization (CDP) extension of the cache allocation interface. When enabled, CDP permits the configuration of separate cache allocation bitmasks for instructions and data.
2.3. Shared Memory Congestion
Although techniques such as caching mitigate the von Neumann bottleneck described in Section 2.2, they do not remove the need to transfer data to and from the main memory. The interconnect between the CPU and the main memory, the memory bus, is limited in bandwidth and shared by all tasks on a non-uniform memory access (NUMA) node. Therefore, memory requests may be required to do additional waiting until sufficient capacity to handle the request is available (pellizzoni_worst_2010). This may result in workloads of low to medium memory intensity becoming drowned out given co-located workloads of high enough memory intensity. An abstract model of the interconnect between the cores, LLC, and main memory is illustrated in Figure 3.
Figure 3. Illustration of the shared memory bus of a modern Intel Xeon Scalable processor (mulnix_intel_2022).
Memory controllers commonly implement a variation of a first-come-first-served (FCFS) request scheduling algorithm that prioritizes row hit requests (rixner_memory_2000). This prioritized queue of requests in the memory controller influences the memory access times of processes (kim_bounding_2016).
2.3.1. Shared Memory Limiting
Partitioning of the memory bus is available on supported hardware, which functionality is provided by a programmable request rate controller implemented in chip hardware. The feature is referred to as memory bandwidth allocation (MBA) by Intel (intel_corporation_intel_2023) and the Linux kernel documentation (yu_24_nodate). However, a more descriptive name is bandwidth limitation as used by AMD (advanced_micro_devices_inc_amd64_2022).
Bandwidth limitation provides the ability to dynamically throttle and delay memory requests from specific processor cores. Typically, a memory-intensive process would have its memory requests limited so that other workloads become less affected by the high bandwidth usage. Bandwidth limits may be specified as a percentage (Intel only), or as a maximum bandwidth in megabytes (yu_24_nodate).
Throttling is applied to all requests that miss L2 cache, the last layer of cache specific to a single physical core. This includes requests that may hit the shared L3 cache (xiang_emba_2019) as illustrated by Figure 4. Thus, bandwidth limitation is the most effective at throttling workloads that are inefficient at utilizing the L3 cache. Throttling is applied at the level of physical cores and any core utilizing simultaneous multi-threading will therefore use the maximum throttling value of its logical cores (herdrich_introduction_2019).
Figure 4. Illustration of memory bandwidth limiting (MBL) impact on the memory request chain. Bandwidth limiting affects all requests that miss the L2 cache, which includes potential L3 hits.
2.4. OS Support for Hardware QoS Enforcement
Management of QoS enforcement schemes is preferably managed through an operating system-level interface. In Linux, this interface is known as the User Interface for Resource Control (resctrl(yu_24_nodate). resctrl enables the definition of resource allocation groups that dictate resource behavior. Logical processor cores or even individual processes may be assigned to a resource group and will subsequently follow the allocation configuration of that group. Resource allocation groups are mapped directly to hardware classes of service (CLOS or COS). Therefore, the number of available resource groups is limited and dependent on the CPU model. When individual processes are assigned to resource groups, the rules of the core executing that process are modified at the time of context switching such that the rules only apply to the correct process.
The resources available to any one process are determined by following a set of rules:
  1. (1)
    If the process is assigned to a resource group, the configuration of that group is used.
  2. (2)
    If the process is in the default resource group but is running on a processor core assigned to a resource group, then the configuration of the core’s group is used.
  3. (3)
    The configuration of the default resource group is used.
2.5. A Note on Shared Resource Monitoring Features
In addition to the enforcement features discussed, the Intel RDT and AMD QoSE feature sets include monitoring support for shared cache and memory bandwidth. These monitoring features are, similarly to their enforcement counterparts, accessible through the resctrl OS interface (yu_24_nodate).
Due to the indirect nature of the mechanisms provided by QoS enforcement features, tuning behavior is highly dependent on the dynamics of the workloads of a system. Thus, actual behavior may be difficult to predict (sohal_closer_2022; farina_assessing_2022; xiang_emba_2019) and it is therefore recommended that actual resource usage is monitored and taken into account by controllers (intel_corporation_intel_2023). Monitoring features therefore play a crucial role in effective feature utilization as they allow controllers to observe changes in metrics resulting from a tuning decision.
3. Review Motivation
Through observation of published literature, it is clear that QoS enforcement hardware features are conspicuous by their absence in general-purpose cloud computing resource management literature reviews, see Table 1.
Table 1. Previous surveys on shared resource management.
Ref. QoS EF* Discussion Context
RT* VNFs* Cloud
(mittal_survey_2017)
(gracioli_survey_2015; lugo_survey_2022)
(kawashima_vision_2021)
(hameed_survey_2016; liaqat_federated_2017; hong_resource_2019; bermejo_virtual_2020; helali_survey_2021; chaurasia_comprehensive_2021; carrion_kubernetes_2022; huang_survey_2023)
This Survey
*Abbreviations — QoS EF: QoS Enforcement Feature Coverage,
RT: Real-time, VNFs: Virtual Network Functions.
Cloud environments rarely see hard execution time deadlines on tasks, as is common in real-time systems, but rather statements of performance intent in the form of service level objectives (SLOs). Resource management optimization for cloud environments is therefore typically discussed in a broader context and at a higher level of abstraction than real-time systems. Thus, lower-level sources of congestion such as LLC and memory bandwidth are often overlooked, excluded, or disregarded in cloud resource management surveys (hameed_survey_2016; liaqat_federated_2017; hong_resource_2019; bermejo_virtual_2020; helali_survey_2021; chaurasia_comprehensive_2021; carrion_kubernetes_2022; huang_survey_2023). Even so, increased resource efficiency and task consolidation achieved through QoS enforcement feature utilization is not exclusive to real-time systems. We thus recognize the need for a review of literature approaching shared resource management by utilizing QoS enforcement features with a general-purpose cloud-computing point of view.
This review serves to gain insight into how the academic community is approaching resource management problems specifically by using hardware-implemented QoS enforcement features, and to identify current problems and opportunities for improvement in general-purpose cloud computing environments.
3.1. Related Surveys
A few systematic reviews have previously been conducted that cover neighboring and overlapping areas of research. In their (gracioli_survey_2015) article, gracioli_survey_2015 (gracioli_survey_2015) surveyed cache management approaches for embedded real-time systems. They recognize that server processors supporting Intel CAT, the only commercially available feature at the time, are not targeted at the embedded systems market surveyed in their paper. However, they still make the case for cache allocation’s usefulness for the optimization of soft real-time applications.
Publications with a focus on resource management in the cloud often approach the problem at a higher abstraction layer than papers with lower-level performance as their context, such as those from the real-time computing community. This often means that CPU cache and memory bandwidth management on the node level is out of scope. Such is the case in the surveys by hameed_survey_2016 (hameed_survey_2016), liaqat_federated_2017 (liaqat_federated_2017), and hong_resource_2019 (hong_resource_2019).
With a clear focus on cache partitioning, mittal_survey_2017 (mittal_survey_2017) surveyed techniques for mitigating the increased shared cache congestion that follows from increasing core counts in processors. This includes early publications utilizing CAT, as well as a large number of software-based approaches using techniques such as page coloring.
bermejo_virtual_2020 (bermejo_virtual_2020), helali_survey_2021 (helali_survey_2021), and chaurasia_comprehensive_2021 (chaurasia_comprehensive_2021) surveyed techniques of workload consolidation in the cloud. The first with a specific focus on virtual machine (VM) consolidation. kawashima_vision_2021 (kawashima_vision_2021) acknowledge the usefulness of cache partitioning when optimizing the performance of virtual network functions (VNFs) that power much of the networking in modern clouds.
lugo_survey_2022 (lugo_survey_2022) surveyed works on process interference for real-time systems published between 2015 and 2020. The works included proposed techniques that reduce congestion in LLC, main memory, as well as in the main memory bus. A few of the works utilize QoS enforcement features, while others use software-based approaches.
With the rise of Kubernetes as the de-facto standard container orchestrator, carrion_kubernetes_2022 (carrion_kubernetes_2022) identifies the need to survey the current approaches and ongoing issues in Kubernetes scheduling. They find that Kubernetes is used in many different environments, running different types of workloads on different types of hardware, which dictates the type of scheduling approaches leveraged.
In a recent publication, huang_survey_2023 (huang_survey_2023) review works on resource management in the context of cloud-native mobile computing. This has received an increasing amount of attention recently due to the architecture of fifth and sixth-generation mobile communication networks (5G and 6G) being based on cloud-native technology. Most of the works reviewed focus on the management of networking resources through the utilization of network slicing or network virtualization.
4. Review Method
This section outlines the systematic approach used when compiling publications leveraging hardware QoS enforcement features. We do this to ensure an unbiased and reproducible selection of publications for inclusion in our review.
4.1. Search Strategy
Our publication search strategy is based on the process outlined by wohlin_guidelines_2014 (wohlin_guidelines_2014) and follows a simple algorithm. First, a starting set of core publications is identified by querying a digital literature database using a defined set of search terms and inclusion criteria. Then, both backward and forward snowballing is used to identify additional papers that fit the inclusion criteria. The snowballing process is repeated after identified papers are added to the set of included papers. This process is repeated until no additional papers that fit the inclusion criteria are found. This process is outlined in Figure 5.
Figure 5. Reference snowballing search strategy used.
Backward snowballing is the process of observing the reference lists of included papers. These lists are scanned and filtered to identify additional papers that meet the inclusion criteria. Similarly, forward snowballing identifies papers that reference already included papers. Any publication that fits the inclusion criteria is added to the set of included papers. The snowballing search space may grow every iteration that one or more papers are added. Once no new paper that fits the inclusion criteria is found during a snowballing iteration, the search is stopped as the set of potential papers no longer grows.
Our motivations for selecting this approach instead of a more traditional database search approach (i.e. kitchenham_evidence-based_2004 (kitchenham_evidence-based_2004)) are as follows: We survey how the academic community leverages a defined set of hardware features in research. We can thus find our starting set of articles by explicitly searching for the names of the features. This should prove a good basis for snowballing as authors typically cite a small set of early papers as sources for further details of a feature. Such publications become natural “roots” in our reference tree. Moreover, it was shown by badampudi_experiences_2015 (badampudi_experiences_2015) that snowballing is comparable, if not more reliable, than a traditional database search approach.
4.2. Threats to Review Validity
The snowballing process is highly dependent on the quality of the starting set of papers used to seed the search. A starting set that is too small, or not diverse enough can leave “dark spots” where many papers are left out. One reason for this is a same author bias that may emerge. As researchers often cite their previous work due to relevance, a snowballing search may become heavily biased towards a few researchers’ works (jalali_systematic_2012). The same phenomenon may be true for publishers instead of individual researchers (wohlin_guidelines_2014).
It is also possible that our review is incomplete as researchers may be using hardware QoS features without explicitly referencing a source for that feature. This would mean that unless the paper is cited by another included paper, it would not be found. There is the possibility that it could be found in the search for the starter set of papers, however, this is unlikely.
4.3. Inclusion Criteria
The criteria used to filter publications are derived from our review objectives. The inclusion criteria used are as follows:
  1. (1)
    The proposal utilizes one or more hardware-based QoS enforcement features available on commercial off-the-shelf x86 chips.
  2. (2)
    The research uses the feature or features first-hand, not through re-implementation of a previous proposal.
  3. (3)
    The publication is four pages or longer.
  4. (4)
    The paper was published before July 1st, 2023.
  5. (5)
    The paper is written in English.
No starting date of the publication time span is defined as hardware QoS features have only recently become available in commodity off-the-shelf server processors. Intel’s CAT was originally introduced in the Xeon E5-2600 v3 family of processors (herdrich_cache_2016) which launched in the third quarter of 2014. Thus, 2014 serves as a starting date for CAT publications.
We limit our review to hardware-based control mechanisms even though software-based QoS enforcement approaches have existed for many years (mittal_survey_2017). Software-based approaches such as page coloring (taylor_tlb_1990) or MemGuard (yun_memguard_2013) provide similar functionality to their hardware-based counterparts and have likely inspired the design of the features now implemented in hardware. However, the native and purpose-built nature of hardware-level control mechanisms means their utilization is preferable in most cases, which has started to make software-based approaches obsolete. Additionally, the implementation specifics of hardware-level control mechanisms mean their areas of applicability and their limitations may vary from software-level approaches.
4.4. Starting Set Identification
Google Scholar [^3] http://scholar.google.com was used as a search engine when querying for potential starting set articles. This had the advantage of minimizing potential publisher bias as the search covers a large number of sources. A separate search was conducted for each of the hardware features of interest. The exact search terms used were:
  1. (1)
    “intel cache allocation technology”
  2. (2)
    “intel memory bandwidth allocation”
  3. (3)
    “amd64 technology platform quality of service extensions”
A set of seed publications were selected from these lists of publications based on our inclusion criteria, publisher diversity, and relevance. This start set of papers are listed in Table 2 and their relationship graph may be seen in Figure 6.
Table 2. Start set of publications used to seed the snowballing search ordered by date of publication from oldest to most recent.
Ref. Search Authors
(funaro_ginseng_2016) 1 funaro_ginseng_2016
(xu_vcat_2017) 1 xu_vcat_2017
(selfa_application_2017) 1 selfa_application_2017
(xu_dcat_2018) 1 xu_dcat_2018
(pons_improving_2018) 1 pons_improving_2018
(park_hypart_2018) 2 park_hypart_2018
(xiang_dcaps_2018) 1 xiang_dcaps_2018
(kim_application_2019) 1 kim_application_2019
(park_copart_2019) 2 park_copart_2019
(xiang_emba_2019) 2 xiang_emba_2019
(zhang_libra_2021) 3 zhang_libra_2021
(sohal_closer_2022) 1 sohal_closer_2022
(farina_assessing_2022) 2 farina_assessing_2022
[…]
[angle=270]./starter_set_svg-tex.pdf_tex
Figure 6. Arc diagram illustrating the reference relationships between the papers selected for the starter set. Publications are chronologically ordered from top to bottom. The size of the nodes corresponds to the number of connections they have.
4.5. Reference Snowballing
Using the snowballing search method, we iterated upon the defined starting set. At the time of the final snowballing algorithm iteration, the search set had increased to $3685$ references. Out of these $3685$ publications, $79$ papers were identified to meet the inclusion criteria of our review. Figure 7 illustrates the included articles’ publication dates and how interest from researchers in both academia and industry has evolved over time.
2015
2016
2017
2018
2019
2020
2021
2022
2023*
0
5
10
15
Papers
Papers using hardware QoS features by year
Academia
Cooperative
Industry
Figure 7. Distribution of articles utilizing platform features grouped by their year of publication. Papers are classified as academic when all authors are affiliated with academic institutions, whereas cooperative publications have one or more authors affiliated with a private company. Industry papers have no affiliation with an academic institution. (*2023 is limited to the first six months of the year.)
It is notable that if reproduced, the search would yield a larger search set than found in this survey. This is due to the included articles being cited as sources in new papers published over time. However, due to the publication date limitation (inclusion criteria (4)), the set of included articles is deterministic and a reproduction will produce the same set of included articles.
5. Taxonomy
We group included publications into three broad categories:
  • Feature evaluations are works that showcase and evaluate the hardware mechanisms from different points of view, in many cases through empirical study.
  • Black-box approaches propose methods or controllers that only observe metrics from hardware or the operating system, treating software as opaque “black boxes”.
  • Application aware approaches observe metrics that are specific to the executing software [^4] E.g. application latencies, throughput, etc. in addition to hardware metrics.
The layout of this section is based on this classification and may be observed in Figure 8.
Figure 8. Taxonomy organization.
5.1. Feature Evaluations
The function of QoS enforcement features implemented in hardware has been experimentally analyzed by a number of different works (sohal_closer_2022; zhang_libra_2021; herdrich_cache_2016; farina_assessing_2022; ihre_sherif_evaluation_2021; schramm_impact_2019; kim_simple_2017; veitch_performance_2017; castiglione_optimal_2018). It is important to verify the behavior of hardware features so that they may be used as a basis for other research. However, the functionality may change between chip generations due to architectural differences (sohal_closer_2022). This further highlights the importance of analyzing the behavior of new implementations as the features mature over time. Additionally, some works (bechtel_denial--service_2022; qiu_is_2021; farshin_reexamining_2020) evaluate if QoS enforcement features may be used as solutions to problems in other research areas. Other publications (shang_coplace_2021) use QoS enforcement features to demonstrate problems and subsequent solutions to resource management problems.
The publications evaluating hardware features are compiled in Table 3. Notably, all of the studies classified as feature evaluations have evaluated and verified the functionality of Intel implementations, AMD QoSE has not been functionally evaluated in the current literature.
Table 3. Publications presenting evaluations of hardware features.
Ref. Evaluation Feature(s)
Intel CAT Intel MBA AMD QoSE
Direct Evaluations
(herdrich_cache_2016) Introduction and feature showcase of CAT.
(sohal_closer_2022) Experimental evaluation of CAT and MBA.
(zhang_libra_2021) Highlights the shortcomings of MBA.
(farina_assessing_2022) Experimental evaluation of MBA for RT systems.
(ihre_sherif_evaluation_2021) Experimental evaluation of CAT.
(schramm_impact_2019) Experimental evaluation of CAT for VNF systems.
(qin_when_2018) Study of cache partitioning in correlation to workload cache characteristics.
(kim_simple_2017) Evaluation of cache partitioning as a tool to meet workload latency targets.
(veitch_performance_2017) Experimental evaluation of CAT for VNF systems.
(castiglione_optimal_2018) Highlights the potential usefulness of CAT for security purposes.
Indirect Evaluations
(qiu_is_2021) Evaluation of FaaS for LC workload hosting.
(bechtel_denial--service_2022) Evaluation of LLC partitioning as DoS protection on iGPU platforms.
(farshin_reexamining_2020) Evaluation of Intel Data Direct I/O.
(shang_coplace_2021) Intelligent Guest to Host Page Mapping.
5.1.1. Direct Evaluations
As a way of introducing Intel CAT to the scientific community, herdrich_cache_2016 (herdrich_cache_2016) present a feature showcase in their (herdrich_cache_2016) paper. The publication addresses the specifics of the implementation and highlights the importance of the feature by suggesting use cases for the new technology. To back up their claims, the effect on the performance of workloads constrained by LLC allocations is presented and discussed.
In their recent study, sohal_closer_2022 (sohal_closer_2022) conduct an experimental evaluation of Intel’s CAT and MBA. It is found that the behavior of the features differs between chip generations. Therefore, they suggest using their evaluation approach for coming architectural versions in order to verify the behavior of such versions.
zhang_libra_2021 (zhang_libra_2021) conducts a smaller study that highlights the shortcomings of Intel MBA. This evaluation is used as motivation when introducing their own software-based approach to memory bandwidth allocation.
Through an experimental analysis, farina_assessing_2022 (farina_assessing_2022) evaluate the effectiveness and behavior of MBA for resource limitation in real-time systems. They conclude that the hardware-based approach to memory bandwidth limiting outperforms similar software-based approaches such as MemGuard (yun_memguard_2013).
In their master thesis, ihre_sherif_evaluation_2021 (ihre_sherif_evaluation_2021) evaluate the performance implications of LLC partitioning on three commonly used workloads: bzip2, redis, and Graph500. They find that tuned partitioning may increase the performance of the aforementioned workloads by up to $5.8\%$, $8.6\%$, and $12.0\%$, respectively, compared to an unmanaged system when co-located with a noisy neighbor.
schramm_impact_2019 (schramm_impact_2019) explore the effectiveness of CAT as a method of improving the performance of virtual network functions (VNFs). They find that sufficient LLC allocation may increase the throughput of a VNF by up to $33\%$ during heavy load scenarios.
To better understand the implications of cache partitioning on workloads with different characteristics, qin_when_2018 (qin_when_2018), kim_simple_2017 (kim_simple_2017), and veitch_performance_2017 (veitch_performance_2017) conduct empirical studies where the behavior of a range of workloads under different cache partitioning schemes is examined. They find that the tuning of cache allocation schemes is generally beneficial to ensuring predictable performance of workloads.
While not strictly presenting a feature evaluation, castiglione_optimal_2018 discuss the potential usefulness of cache allocation in preventing and mitigating cache side-channel attacks (castiglione_optimal_2018). They observe the behavior enabling the side-channel attack vector and argue for the applicability of cache partitioning as a possible solution to the problem. However, determining if cache allocation upholds the promises is left as future work.
5.1.2. Indirect Evaluations
With the rise of serverless computing and function-as-a-service (FaaS) platforms, an increasing number of practitioners are looking to the promises of ease-of-use and pay-per-use of the paradigm. qiu_is_2021 (qiu_is_2021) recognize the need to evaluate the feasibility of FaaS as a model to host LC workloads. In their evaluation, they include LLC and memory bandwidth partitioning as potential ways of ensuring consistent performance. Their findings conclude that FaaS is not yet ready for LC workloads and that the aforementioned partitioning techniques would have to support orders of magnitude more CLOSs before QoS of FaaS functions can be sufficiently enforced.
On platforms with integrated graphics, LLC is often shared between the CPU and GPU. This becomes an attack vector for denial-of-service (DoS) attacks. bechtel_denial--service_2022 (bechtel_denial--service_2022) evaluates if LLC partitioning using CAT provides sufficient protection to mitigate such attacks. They conclude that CAT may indeed provide spatial isolation, but that there are still cases where temporal isolation can not be achieved. This in turn leads to sophisticated attacks still being viable.
To better understand the mechanics of Intel’s Data Direct I/O (DDIO) and its correlation to LLC cache ways, farshin_reexamining_2020 (farshin_reexamining_2020) conducts an empirical study to gain a better understanding of the inner workings of the technology. With the knowledge gained from this study, they discuss its effectiveness and shortcomings. General advice regarding how DDIO should be utilized depending on the use case is also presented.
shang_coplace_2021 (shang_coplace_2021) utilize LLC partitioning to demonstrate that caching inefficiencies exist in the way hosts map guest VM memory pages to physical memory. Their proposed solution, CoPlace mitigates this problem and reduces cache conflict misses that exist even when a guest is given exclusive access to some subset of LLC.
5.1.3. Summary
The hardware features of interest have been evaluated in many works from different points of view. Some with the intention of showcasing and introducing them to the scientific community, and some to see if they can help solve a problem in another research field. It is clear that the behavior may change between chip architecture generations, especially in more recent and less mature implementations. Behavior should thus be validated in future generations as the implementations mature.
5.2. Black-Box Approaches
The ability to tune a system without knowledge of the applications it is hosting can be useful. However, doing so may be challenging considering that different applications are sensitive to different bottlenecks (cortez_resource_2017). Even so, it may be required in some cases due to privacy concerns. Such is the case with public cloud providers that are unable to observe the workloads running in customers’ virtual machines (shahrad_provisioning_2021). Consequently, tuning must be done by observation of system metrics only. Thus, tuning becomes workload agnostic, as all processes are observed using the same system and operating system metrics.
Proposals leveraging the surveyed platform features that treat workloads as “black boxes” are presented in Table 4. We split black box approaches into a number of subclasses: Proactive approaches, reactive approaches, combined approaches, workload clustering approaches, and feature abstractions.
Table 4. Publications leveraging only hardware and operating system-level metrics.
Ref. CA* MBL* DC* Optimization Proposed Technique
Proactive Approaches
(xiao_cppf_2019) Generic Performance Prefetch-Aware Heuristic LLC Partitioning
(kim_application_2019) Generic Perf. Prediction ML Performance Prediction Model
(aupy_co-scheduling_2019) HPC Perf. Prediction Mathematical Modelling of Execution Times
(durner_towards_2019) VNFs Utilization Static Heuristic Optimization Algorithm
(funaro_ginseng_2016) Cloud Performance Market-Driven Auction-Based Scheme
(shahrad_provisioning_2021) Cloud Performance Cache-Aware VM Scheduler
(liu_catalyst_2016) Security Security LLC Side-channel Attack Protection Scheme
(dessouky_chunked-cache_2021) Security Perf., Security LLC Side-channel Attack Protection Scheme
(zeng_playing_2021) Generic Performance Greedy Static LLC Allocation Scheme
(chung_enforcing_2019) Generic Performance LLC-Aware Memory Bandwidth Management
(thomas_dark_2018) VNFs Performance Memory Hierarchy Management Framework
(sprabery_novel_2017) Security Security LLC Side-Channel Attack Protection Scheme
Reactive Approaches
(navarro-torres_balancer_2023) Generic Perf., Fairness Heuristic LLC and MB Management Scheme
(park_hypart_2018) Generic Performance Explorative MB Partitioning Algorithm
(park_copart_2019) Generic Perf., Fairness Explorative LLC and MB Partitioning
(xu_dcat_2018) Generic Performance Heuristic Program Phase Detection
(pons_cache-poll_2022) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(farina_enabling_2023) RT Performance Heuristic Bandwidth Guarantees
(roy_satori_2021) Generic Perf., Fairness Bayesian Optimization Based Algorithm
(chen_alita_2020) Cloud Fairness Heuristic VM Resource Fairness Algorithm
(xiang_dcaps_2018) Generic Performance Simulated Annealing Optimization
(yuan_dont_2021) Generic Performance Heuristic, DDIO Aware LLC Management
(yi_mt2_2022) Generic Performance Heuristic MB Management Algorithm
(yao_leveraging_2019) Security Security Cache Timing Channel Attack Mitigation
Combined Approaches
(li_pfa_2022) Generic Perf., Fairness Heuristic LLC Management Algorithm
(garcia-garcia_lfoc_2019) Generic Fairness Heuristic Cache-Clustering Algorithm
(saez_lfoc_2022) Generic Fairness Heuristic Cache-Clustering Algorithm
(pons_improving_2018) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(pons_phase-aware_2020) Generic Perf., Fairness Heuristic LLC Partitioning Scheme
(chen_drlpart_2021) Generic Performance Deep Reinforcement Learning Model
(chatterjee_com-cas_2023) Generic Performance Compiler Guided LLC Management
(zhang_zeus_2021) Cloud Utilization Kubernetes Resource Management Scheme
(zhu_kelp_2019) ML Performance Data Management for ML Systems
Workload Clustering Approaches
(tang_themis_2023) Generic Perf., Fairness K-Means Classification of Guest VMs
(gifford_dna_2021) Generic Performance Workload Clustering Technique
(xiang_emba_2019) Generic Performance Heuristic Hierarchical Clustering Algorithm
(selfa_application_2017) Generic Fairness K-Means Classification of Workloads
(qiu_machine-learning-based_2022) Cloud Performance SVM Classification and BO LLC Partitioning
Feature Abstractions
(xu_vcat_2017) VMs Performance Virtualization of CAT
*Abbreviations — CA: Cache Allocation, MBL: Memory Bandwidth Limiting, DC: Discussion Context.
5.2.1. Proactive Approaches
Publications featuring approaches classified as proactive attempts to make resource management decisions ahead of time. They may enforce static partitioning schemes or try to predict future resource requirements of workloads.
By profiling workloads as sensitive, or non-sensitive, to hardware prefetching of memory, xiao_cppf_2019 (xiao_cppf_2019) propose a controller that aims to improve LLC management on modern systems equipped with hardware prefetching. They call their approach CPpf, which partitions LLC in accordance with the determined workload classifications. The system allocates a single exclusive cache way to each prefetching sensitive workload, leaving the rest to be shared among non-prefetching sensitive workloads.
With the goal of predicting how a process’ instructions per cycle (IPC) metric changes as a result of a change in its available LLC, kim_application_2019 (kim_application_2019) propose POCAT. POCAT is a trained machine learning model that can predict changes to IPC with an error of 4.7% without affecting current LLC allocations.
To get a better understanding of how runtimes of BE high-performance computing (HPC) workloads are affected by core and LLC allocations, aupy_co-scheduling_2019 (aupy_co-scheduling_2019) create a mathematical model that predicts how job execution times are affected by changes in scarce shared resource availability. Using their model, they are able to more efficiently partition resources among workloads to increase the efficiency of a system.
A simple heuristic algorithm to partition LLC among VNFs is proposed by durner_towards_2019 (durner_towards_2019). Their algorithm distributes available cache ways based on CPU utilization, which results in lowered total system utilization, increasing energy efficiency and QoS. However, only static configurations are considered where virtual network functions (VNFs) and their loads do not change. This makes their algorithm valid for very specific use cases only.
In an attempt to make LLC partitioning more effective, Ginseng (funaro_ginseng_2016) allow consumers (i.e. workloads) to bid on LLC space using Vickrey-Clarke-Groves auctioning. This allocates LLC to the process that would see the largest economic benefit from having said space. As the auctioning and allocation cycle is kept short, processes may acquire more LLC space at critical phases in its operation, while seceding it at times of less sensitivity.
Public cloud providers aim to reduce their total cost of hardware ownership by optimizing its utilization. Making a VM scheduler LLC aware allows the provider to give LLC QoS guarantees to their customers as LLC ways may be exclusively, or partially exclusively, allocated to specific VMs. This is the goal of CacheSlicer, a VM scheduler proposed by shahrad_provisioning_2021 (shahrad_provisioning_2021) that schedules VMs onto nodes based on low-cost, normal, and cache-sensitive VM categories.
In public clouds, where different clients’ VMs share physical hosts, malicious actors may compromise security by extracting sensitive information from a co-located guest by manipulation of the LLC. CATalyst, as proposed by liu_catalyst_2016 (liu_catalyst_2016), is a module that integrates into hypervisors that protect against such LLC side-channel attacks by partitioning the LLC to provide security guarantees.
To improve general performance in trusted execution environments, dessouky_chunked-cache_2021 propose a hybrid cache management scheme that splits applications based on their sensitivity to LLC side-channel attacks. Their approach, titled Chunked-Cache (dessouky_chunked-cache_2021), enforces cache security for sensitive applications by placing them in an isolated domain. One such domain is created and enforced for each sensitive context, while the operating system and other insensitive workloads share a non-sensitive domain which improves overall system performance to similar side-channel protection schemes.
zeng_playing_2021 find that a simple static allocation of LLC is not sufficient to improve system performance when software prefetching of memory is used (zeng_playing_2021). They show that improvements can be made even with a simple greedy partitioning scheme.
An observation that allocating more LLC to an LC application may actually bring worse performance is made by chung_enforcing_2019 (chung_enforcing_2019). This is due to the memory controller being overwhelmed by the requests of the remaining workloads that miss LLC more often due to less remaining usable after partitioning. The paper details their approach to mitigate this issue and the “memory virtual channels” they use to prioritize and order memory requests such that the LC application is guaranteed prioritized access to main memory as well as LLC.
thomas_dark_2018 observe that as networks approach transfer speeds of 100 gigabits per second and beyond, the memory system becomes a bottleneck that can a cause large number of packet drops if incorrectly utilized (thomas_dark_2018). As an attempt to mitigate this issue, they propose CacheBuilder. Cache builder is a system that, among other techniques, utilizes cache allocation to provide applications with limited control of the memory hierarchy such that packet drops may be minimized.
By splitting LLC into two partitions, sprabery_novel_2017 (sprabery_novel_2017) attempts to protect sensitive applications against cache side-channel attacks. Sensitive workloads are assigned to one, while other, insensitive, workloads share the other. The cores assigned to the sensitive partition will upon context switching perform a “state cleansing”. This means that the LLC used by the previous execution context will be flushed before the new context begins execution. This approach protects sensitive data at the cost of performance.
5.2.2. Reactive Approaches
By the continuous observation of system metrics, reactive approaches may adjust resource partitioning in reaction to changing system state.
navarro-torres_balancer_2023 (navarro-torres_balancer_2023) identify changes in program phase by observing cache and memory bandwidth usage behaviors. Such changes trigger reevaluations of the current allocation scheme for LLC and memory bandwidth, respectively. The two resource dimensions are controlled by separate heuristic schemes, that work together to balance resource congestion. This controller is called Balancer, which aims to improve overall system performance and thread fairness.
HyPart (park_hypart_2018) is a dynamic memory bandwidth partitioning controller with a focus on fairness proposed by park_hypart_2018 that uses workload profiling as a base for its explorative approach to resource partitioning. Profiling workload behavior allows the scheme to eliminate potential states from the state space it uses to iteratively and heuristically converge on some optimum partitioning scheme. The process is repeated should the state of the system change. By expanding upon this work, the authors later propose CoPart (park_copart_2019). In this extended work, the controller is able to optimize for performance in addition to fairness. It is also extended to partition and manage LLC allocation instead of solely managing memory bandwidth.
xu_dcat_2018 (xu_dcat_2018) propose dCat, a dynamic LLC partitioning controller that aims to improve the performance of system workloads. This is done through heuristic workload phase detection based on cache access, instruction retirement, and instructions per cycle (IPC) metrics. Workloads are classified into the classes reclaim, receiver, donor, keeper, streaming, and unknown based on their current cache utilization behavior. Allocation changes are triggered by applications transitioning between states due to changes in cache access behavior.
pons_cache-poll_2022 (pons_cache-poll_2022) aim to improve system fairness and performance by identifying cache-polluting processes and limiting their ability to impact cache-sensitive applications with Cache-Poll. Their approach is a heuristic algorithm that classifies applications based on system metrics into one of three classes. Cache sensitive, mildly cache sensitive, and cache insensitive. Based on the said classification, Cache-Poll then partitions available LLC in order to improve performance and fairness.
By building upon their work presented in (farina_assessing_2022), farina_enabling_2023 (farina_enabling_2023) propose a heuristic for estimating and guaranteeing memory access for critical cores while not significantly sacrificing the performance of non-critical ones. Their approach is based on observation and analysis of the un-core metric read pending queue (RPQ) found on Intel processors.
roy_satori_2021 roy_satori_2021 increase system throughput and fairness of multiple BE jobs competing over system resources. They do this by proposing Satori, a resource partitioning controller utilizing Bayesian optimization to optimize cache and memory bandwidth allocations based on a novel objective function using system metrics.
By observing a tenant VM’s ability to significantly reduce the performance of co-located VMs, chen_alita_2020 (chen_alita_2020) argue that a dynamic resource management system is needed to improve fairness among tenants. Their proposal called Alita identifies resource polluters and limits their availability to resources that are under pressure. The identification of such polluters is done solely through the use of hardware and OS-level metrics.
xiang_dcaps_2018 (xiang_dcaps_2018) approach the problem of cache partitioning by searching the large space of possible partitioning combinations for a near-optimal configuration. They do this by proposing an algorithm based on the simulated annealing optimization scheme that optimizes workload performance. This scheme is called dynamic allocation with partial sharing (DCAPS), and operates in user space, using only system metrics.
Intel’s Data Direct I/O (DDIO) technology enables the transfer of data directly between network controllers and LLC, completely bypassing main memory and increasing performance. However, this may introduce additional LLC congestion. In an attempt to mitigate this, yuan_dont_2021 (yuan_dont_2021) propose IAT. IAT is a heuristic cache partitioning scheme that resembles a Mealy finite state machine (FSM). Workloads are classified into states that map onto states of the FSM. Based on system metrics, workloads may change state and thus trigger LLC reallocation. IAT is DDIO aware and may allocate part of the LLC exclusively to DDIO to improve system performance.
Through observation of system metrics, yi_mt2_2022 (yi_mt2_2022) aim to improve system performance on platforms with non-volatile memory (NVM) by regulating memory bandwidth with their MT2 controller. The kernel-space controller allows administrator control of its throttling groups, which are based on the Linux kernel’s cgroups feature. By continuously updating the bandwidth throttling settings based on application bandwidth usage, overall system performance may be improved.
To protect against cache timing channel attacks between hardware threads, applications, or VMs, yao_leveraging_2019 proposes a system that monitors cache behavior and heuristically identifies pairs of workloads with suspiciously correlated activity (yao_leveraging_2019). Any suspected spy-victim pair are separated by strictly separating their available cache.
5.2.3. Combined Approaches
By leveraging both proactive and reactive strategies, combined approaches are able to create approximate partitioning schemes that may be fine-tuned by observing system metrics.
By classifying workloads based on cache usage behavior, li_pfa_2022 (li_pfa_2022) propose a performance and fairness-aware LLC partitioning scheme called PFA. Classification is performed through the observation of cache hit rates of retired CPU instructions, after which a prediction of LLC requirements is made. During operation, performance slack is monitored and partitioning is adjusted accordingly. Improvements are observed in both performance and fairness using their controller, but it is noted that further improvements should be achievable by tuning memory bandwidth in combination with LLC partitioning.
With a clear focus on fairness, garcia-garcia_lfoc_2019 (garcia-garcia_lfoc_2019) propose LFOC, a controller that classifies workloads by their cache behavior. The classes defined are light sharing, streaming, and sensitive. LLC is then partitioned accordingly using a heuristic cache-clustering algorithm to increase overall system fairness. This work is later extended in LFOC+ (saez_lfoc_2022) where an improved pair-clustering strategy is used to further improve system fairness.
By dividing the LLC into two partitions, pons_improving_2018 (pons_improving_2018) attempt to improve system performance and fairness. Workloads are profiled offline and classified as either cache sensitive or cache insensitive based on how they respond to changes in cache availability, after which they are assigned to the corresponding LLC partition. By dynamically “moving the goalpost”, i.e. shrinking one partition and increasing the other, they aim to decrease system congestion. Their heuristic approach is called critical-aware or simply referred to as CA. This work is later extended in CPA (pons_phase-aware_2020) which uses instructions per cycle and cache hit rate metrics to determine the phase of a workload and allocates LLC accordingly. Here, LLC partitioning is not limited to just two partitions.
Through their use of deep reinforcement learning (DRL), chen_drlpart_2021 (chen_drlpart_2021) present a resource allocation controller called DRLPart that partitions the resources of a system among workloads using a reward prediction model. The goal of the allocations is to reduce resource congestion and increase workload throughput. To validate its effectiveness, a comparison with two older controllers (patel_clite_2020; xiang_dcaps_2018) is carried out.
By allowing the compiler to create hints about the cache behavior of programs, chatterjee_com-cas_2023 (chatterjee_com-cas_2023) are able to make proactive, just-in-time, LLC allocation decisions. Their proposed scheme, Com-CAS, makes use of a compilation component based on LLVM and a runtime component that works in unison to achieve improved system performance.
With the goal of mitigating resource under-utilization in large-scale Kubernetes clusters, zhang_zeus_2021 detail their resource management scheme Zeus (zhang_zeus_2021). Zeus intercepts Kubernetes API requests and modifies allocation requests in accordance with its observed node-level metrics. On each node, LLC is split into two partitions. One for LC workloads, and one for BE jobs. This allows the controller to minimize interference and thus increase average, cluster-wide, CPU utilization from $15\%$ to $60\%$.
To improve the performance of accelerated ML workloads, zhu_kelp_2019 leverage non-uniform memory access (NUMA) node subdomains in their Kelp (zhu_kelp_2019) runtime. Each system socket is split into two subdomains, one for high-priority ML workloads and one for low-priority tasks. This separation in combination with heuristic resource partitioning reduces task interference such that ML workload throughput is increased.
5.2.4. Workload Clustering Approaches
Approaches that cluster workloads attempt to group workloads by behavior so that allocation of resources may be performed in a generalized fashion.
Themis (tang_themis_2023) by tang_themis_2023 increases QoS and fairness of guest VMs by observing their memory bandwidth usage behavior. The VMs are classified into groups using the Euclidean K-Means algorithm, where the number of groups equal the number of available COS on the system. The controller then proactively partitions LLC and MB by following a heuristic algorithm. The classification and partitioning scheme is reevaluated over time by continuously observing VM bandwidth usage. The controller may also react to saturation of the memory bus by moving VMs to other NUMA nodes if available.
By observing a core’s LLC cache miss percentage and number of retired instructions, gifford_dna_2021 (gifford_dna_2021) classify workloads into clusters depending on the program phase they are deemed to currently be in. These clusters are then allocated LLC ways and memory bandwidth555 Memory bandwidth is allocated using MemGuard (yun_memguard_2013), a software-based approach. in order to maximize the performance of a system based on a heuristic algorithm. They call their approach DNA. A deadline-aware version that is able to take a real-time workload’s execution deadlines into consideration called DADNA is also presented. Their experimental evaluation of the controller only compares DNA and DADNA against (xu_holistic_2019), another controller previously proposed by the same authors.
With the goal of increasing overall system performance, xiang_emba_2019 (xiang_emba_2019) propose EMBA. EMBA is a controller that clusters workloads using a heuristic hierarchical clustering algorithm based on their usage of the memory bus. This classification is then used to heuristically and iteratively achieve a memory bandwidth throttle scheme that improves overall performance.
With system fairness as their goal, selfa_application_2017 (selfa_application_2017) propose a set of LLC allocation policies that are based on workload clustering using the K-Means algorithm (hartigan_algorithm_1979). The classification is done based on the number of stalled CPU cycles due to L2 cache misses. Then, the LLC is partitioned accordingly which results in an improvement of overall system fairness.
By clustering workloads using a support vector machine (SVM) based machine learning model, qiu_machine-learning-based_2022 are able to classify workload behavior using only system metrics. This is the first phase of their Clustering-and-Allocation (C&A) (qiu_machine-learning-based_2022) LLC partitioning scheme. Using the aforementioned workload classification as input, a Bayesian optimization-based approach is used to quickly and efficiently reach a near-optimum partitioning scheme. In their own evaluation, C&A is found to increase throughput by up to $26.2\%$ in a best-case scenario.
5.2.5. Feature Abstractions
Due to the nature of hardware features such as Intel CAT, they are limited to bare-metal deployments with direct access to hardware. In an attempt to bring LLC allocation to virtualized environments, xu_vcat_2017 (xu_vcat_2017) propose vCAT. vCAT is a system that exposes virtual cache partitioning to VMs, which is later mapped to physical LLC partitions on the host. By the extension of the functionality of hypervisors, guests are able to give hints regarding their workloads’ requirements, while the host remains in control.
5.2.6. Summary
Hardware QoS enforcement features may be used to great effect and tuned in many ways even when only system metrics are observed. Throughput may be increased by $26.2\%$ (qiu_machine-learning-based_2022) or $18.5\%$ (chen_drlpart_2021), performance increased by $7.1\%$ (navarro-torres_balancer_2023), or fairness improved by $44\%$ (pons_cache-poll_2022) compared to an unmanaged system.
Many publications classify workloads with regard to their utilization and sensitivity to a shortage of shared system resources. This categorization is then often used to dictate how resources are partitioned among workloads. Other papers try to predict how changes in resource allocation would impact the performance of workloads through machine learning and mathematical approaches. Conversely, reactive controllers have been proposed that continuously profile the workloads and reactively adjust resource allocations as requirements change over time. Additionally, publications focused on security (liu_catalyst_2016), compiler inserted cache behavior hints (chatterjee_com-cas_2023), auctioning-based partitioning (funaro_ginseng_2016), and LLC management on the data-center scale (shahrad_provisioning_2021), and more have been discussed.
5.3. Application Aware Approaches
To better understand how shared resource allocation affects the real-world performance of workloads, controllers may monitor key performance indicators (KPIs) of applications. Such metrics may include request throughput, average latencies, tail latencies, and more. Every workload’s KPIs are different and manual configuration of metric probes is therefore often required. Resource management controllers need to be aware of how to measure an application’s metrics, as well as their target values and the importance of one metric compared to another. Many workloads have service-level objectives (SLOs) to adhere to which allow application-aware resource managers to minimize the resource availability of such workloads while ensuring they still operate within their targets. Thus, freeing more resources to maximize the performance of other co-located workloads.
We group application-aware approaches into four categories: Proactive, reactive, combined, and performance slack management approaches. All identified papers that fit the application-aware category are listed in Table 5.
Table 5. Publications leveraging both system and application metrics.
Ref. CA* MBL* DC* Optimization Proposed Technique
Proactive Approaches
(jeatsa_casy_2022) Cloud Performance ML Resource Management Controller for FaaS
(stewart_performance_2022) Generic Performance ML Short-term LLC Allocation Controller
(noll_accelerating_2018) DBMS Performance LLC Management for DBMS
Reactive Approaches
(nikas_dicer_2019) Generic Utilization Heuristic LLC Partitioning Scheme
(chintapalli_restrain_2022) VNFs Performance Heuristic VNF Resource Management
(duan_improving_2021) Generic Performance ML Classification and Heuristic Scheduling
(duan_improving_2022) Generic Performance ML Classification and Heuristic Scheduling
(patel_clite_2020) Cloud Performance Bayesian Optimization Based Algorithm
(pons_effect_2022) Generic Performance Analysis of App Sensitivity to Shared Resources
(gureya_generalizing_2021) Cloud Performance Heuristic Multi-Socket Memory Management
(papadakis_improving_2017) Generic Performance Heuristic LLC Partitioning Scheme
(qin_locpart_2017) Generic Performance Heuristic LLC Partitioning Scheme
Combined Approaches
(qiu_slo_2021) Generic Performance Profiling & Feedback Based Management
(nejat_coordinated_2020) Generic Energy Efficiency Modeling of Workload Behavior
(zhao_rhythm_2020) Cloud Utilization Sub-service LLC Management Scheme
(tootoonchian_resq_2018) VNFs Utilization Greedy Profiling Based LLC Allocation
(metsch_intent-driven_2023) Cloud Performance ML and Intent-driven Resource Management
(li_rldrm_2020) VNFs Utilization Deep Reinforcement Learning LLC Management
(zhang_libra_2021) Generic Performance Heuristic Memory Bandwidth Management
(chintapalli_ravin_2023) VNFs Utilization Heuristic VNF and Shared Resource Allocation
(xu_holistic_2019) RT Performance Heuristic LLC and BW Management Algorithm
Performance Slack Management Approaches
(zhu_dirigent_2016) Generic Utilization Execution Time Prediction Algorithm
(chen_parties_2019) Cloud Utilization Heuristic Slack Management Algorithm
(lo_heracles_2015) Cloud Performance Heuristic Slack Management Algorithm
(nejat_cooperative_2022) Generic Energy Efficiency Heuristic Slack Management Algorithm
(pang_adaptive_2021) Cloud Energy Efficiency Multi-level Resource Management Scheme
(chen_avalon_2019) Generic Utilization Runtime Prediction and LLC Management
*Abbreviations — CA: Cache Allocation, MBL: Memory Bandwidth Limiting, DC: Discussion Context.
5.3.1. Proactive Approaches
By predicting the resource requirements of workloads (jeatsa_casy_2022; stewart_performance_2022), or by observing workload characteristics to identify beneficial static configurations (noll_accelerating_2018), proactive approaches improve system goals ahead of time.
CASY (jeatsa_casy_2022) by jeatsa_casy_2022 is a cache allocation controller for function-as-a-service (FaaS) systems that uses machine learning to build a cache usage profile of functions with regard to their input values. This allows the controller to tune functions’ allocation of cache by examining the incoming input and thus improve system-wide performance.
With the use of deep learning, stewart_performance_2022 (stewart_performance_2022) profile web services’ cache usage characteristics. This allows them to propose their short-term cache allocation controller that observes response times of web services and temporarily allocates additional LLC to particularly slow requests in order to minimize SLO infringement.
Database management systems (DBMS) are required to handle large amounts of queries and data, but all operations are not equal in cache sensitivity and efficiency as noted by noll_accelerating_2018 (noll_accelerating_2018). They find that certain memory-intensive operations such as a column scan are insensitive to available cache, while simultaneously thrashing existing cache contents. Thus, they propose allocating the minimum possible cache to worker threads that handle such queries which reduces overall cache pollution and thus increases system performance by, in their evaluation, up to $38\%$.
5.3.2. Reactive Approaches
Application-aware reactive approaches monitor application metrics and identify required adjustments to resource partitioning schemes as application load and type of load change over time. They react by applying corrective allocation adjustments that allow optimization of the controllers’ respective goals.
Through observation of both application metrics and system metrics, DICER (nikas_dicer_2019) classifies workloads based on their sensitivity to LLC availability. Additionally, it monitors MB saturation and detects workload phase changes by identifying changes in LLC utilization. The controller proposed by nikas_dicer_2019 uses the gathered data to heuristically increase system utilization by improving BE job performance without allowing LC workloads to break QoS targets.
Virtual network functions (VNFs) often have strict performance guarantees to meet. chintapalli_restrain_2022 (chintapalli_restrain_2022) argue that to be able to co-locate multiple VNFs with other services while ensuring QoS, isolation of resources is crucial. They present RESTRAIN, a dynamic LLC and MB allocation scheme that continuously classifies VNFs into one of four categories using a modified technique borrowed from park_copart_2019 (park_copart_2019). The categories used are “donor”, “receiver”, “donor and receiver”, and “none”. A donor is a workload that is able to release a certain resource and still meet QoS, while receivers need more resources to meet QoS. Donors and receivers are workloads that require more of one resource while being in excess of another. Workloads that are at a resource allocation are classified as none. The controller builds upon and is compared to ResQ (tootoonchian_resq_2018).
With large scale data-centers in mind, duan_improving_2021 (duan_improving_2021) propose a machine learning-based workload classification model. This model is later used by a local search server re-consolidation algorithm. The controller utilizes LLC partitioning and core allocation to guarantee workload performance during re-consolidation from overloaded machines. This work is continued and extended by the same authors in duan_improving_2022 (duan_improving_2022).
With the goal of increasing system utilization by co-locating a greater number of BE jobs without breaking LC service SLOs, patel_clite_2020 (patel_clite_2020) propose CLITE. CLITE is a memory bandwidth and LLC partitioning controller that uses Bayesian optimization to find an optimal configuration of the system. In their case, the controller maximizes the performance of BE jobs while LC applications still meet their QoS requirements.
Identifying the most critical resource of a workload at runtime can be difficult. pons_effect_2022 (pons_effect_2022) try to show how different types of workloads behave when constrained by different kinds of resources. This gives an insight into how resource allocation features behave when used in conjunction. For example, they show that a reduction in LLC availability may increase the memory bandwidth used by a process. This allows them to present a number of key takeaways and guidelines that may help guide cloud operators to make more efficient use of their hardware by reducing resource allocation to insensitive workloads.
gureya_generalizing_2021 (gureya_generalizing_2021) recognize that Intel’s MBA is limited to single socket control only due to the details of its implementation. In an attempt to bring holistic memory bandwidth allocation to multi-socket systems, they propose BALM. BALM is a controller that combines a cross-socket page migration scheme with knowledge of workloads’ SLO targets. This allows the controller to achieve better overall system performance through memory bandwidth partitioning and socket migration of workloads while maintaining SLO targets.
By allocating an exclusive cache partition to a high-priority workload while other applications share the rest, the QoS of the high-priority workload can be maintained. This is what papadakis_improving_2017 propose with their heuristic Dynamic Cache-Partitioning QoS scheme (papadakis_improving_2017). The controller observes the performance of the high-priority workload and adjusts its LLC allocation such that QoS is maintained while maximizing LLC available to remaining system workloads.
qin_locpart_2017 approach LLC partitioning by allocating each workload a separate exclusive piece of LLC in Locpart (qin_locpart_2017). By monitoring the QoS of LC workloads, LLC may be redistributed to or from LC workloads to BE jobs in such a way that LC services meet their performance targets and BE job performance is maximized. A limitation of the approach is that the number of available COSs dictates the maximum number of total workloads assigned to a system.
5.3.3. Combined Approaches
Combined application-aware approaches utilize a combination of proactive and reactive techniques to achieve their goals. Intuitively, this means they perform some ahead-of-time prediction of resource requirements, perhaps through offline application profiling. Then, at runtime, the workloads are monitored and minor corrective adjustments are performed as load and workload behavior changes over time.
With the goal of increasing system performance while maintaining LC workload SLO requirements, qiu_slo_2021 (qiu_slo_2021) propose CoCo. CoCo is an SLO-aware LLC and memory bandwidth partitioning controller that profiles workloads’ sensitivity to such scarce resources. With this knowledge, it dynamically allocates resources based on available SLO overhead to increase overall system utilization and performance.
nejat_coordinated_2020 have total system efficiency in mind when proposing their resource management controller (nejat_coordinated_2020). By observing the SLOs and metrics of workloads, their controller is able to reduce power consumption while still meeting those QoS requirements. This is achieved by modeling the cache and power usage behavior of workloads and then partitioning LLC and scaling core voltages to reduce energy consumption while still allowing applications to meet their intended targets.
By gathering multiple co-located parts of the same LC service into a new abstraction called a Servpod, zhao_rhythm_2020 (zhao_rhythm_2020) are able to improve system utilization while ensuring that LC workloads meet their SLOs. This is achieved by offline profiling and classification of Servpods based on their contributions to overall system tail-latency and sensitivity to shortage of shared scarce resources. They call their scheme Rhythm, which partitions LLC, manages core voltages, etc., with the goal of increasing utilization without LC SLO breakage.
By using the result from offline profiling of VNFs prior to deployment, ResQ (tootoonchian_resq_2018) is able to make greedy decisions regarding LLC allocation optimizations. The controller, proposed by tootoonchian_resq_2018, ensures VNF QoS requirements are met while maximizing the performance of co-located BE workloads.
To allow workload owners to specify SLO targets instead of resource allocation requests in the cloud, metsch_intent-driven_2023 (metsch_intent-driven_2023) propose their intent driven orchestrator. The system learns from application behavior and how changes in the allocation of resources such as LLC affect SLO metrics. This allows the controller to predict how changes in resource allocation will affect application performance and make resource partitioning decisions accordingly.
Due to strict SLO requirements, VNFs are often placed on underutilized machines so that their performance can be guaranteed. li_rldrm_2020 (li_rldrm_2020) propose RLDRM in an attempt to improve utilization of such machines. RLDRM is a closed-loop deep reinforcement learning-based LLC partitioning controller. By learning from the behavior of workloads the controller can make sure the high-priority VNFs have enough LLC ways to meet their SLO requirements while allowing a greater number of BE jobs to consume the remaining resources.
After identifying the shortcomings of Intel’s MBA, zhang_libra_2021 (zhang_libra_2021) propose a custom memory bandwidth partitioning mechanism they call LIBRA. By custom manipulation of the read pending queue, finer-grained control is achieved than what is available through MBA. LIBRA still leverages CAT as a mechanism of LLC partitioning.
Through prior profiling of VNF sensitivity to cache and memory bandwidth constraints, chintapalli_ravin_2023 (chintapalli_ravin_2023) propose RAVIN. The heuristic VNF orchestrator schedules VNFs across machines while considering available cache and memory bandwidth such that throughput guarantees may be met. This is done using the previously gathered profiling information which provides the orchestrator with minimum resource requirements depending on desired throughput guarantees. This allows RAVIN to efficiently utilize available hardware through more efficient packing of VNFs.
To ensure that the worst-case execution time (WCET) of real-time workloads stays within requirements, xu_holistic_2019 propose CaM (xu_holistic_2019). The three-phase heuristic algorithm groups workloads based on their sensitivity to resource shortage before it load-balances them across the cores of a system. LLC and memory BW (using MemGuard (yun_memguard_2013)) are then allocated such that WCET targets are fulfilled.
5.3.4. Performance Slack Management Approaches
Management of performance slack is a concept commonly seen in the context of real-time systems. When operating under completion deadlines, workloads may be tuned such that they may operate “ahead of schedule” which allows performance to be prioritized elsewhere at times of need. This notion of slack is used by several application-aware approaches utilizing hardware QoS enforcement.
With knowledge of an LC workload’s QoS requirements and prior offline profiling, Dirigent (zhu_dirigent_2016) may manage LLC in such a way that co-located BE jobs’ performance can be maximized and thus increase overall system utilization. zhu_dirigent_2016 approach the problem by predicting the remaining runtime of workloads using data gathered during the prior profiling. By comparing how it correlates to set deadlines, resources may be adjusted dynamically so that the performance of LC workloads meets set targets while allowing an increased amount of resources to co-located jobs.
With the premise that only a single LC workload is located on any one machine to guarantee QoS, chen_parties_2019 (chen_parties_2019) propose PARTIES. PARTIES is a QoS-aware resource partitioning controller that enables the co-location and QoS enforcement of multiple LC workloads while maximizing the performance of BE jobs to improve system utilization. This is achieved by observing the performance slack of LC workloads and increasing resource allocation of services about to break their targets.
In one of the first works utilizing platform features, lo_heracles_2015 propose Heracles (lo_heracles_2015), a resource partitioning controller that monitors QoS metrics of an LC workload. By observing that LC workload’s performance slack, its resource allocations may be adjusted such that just enough resources are available to meet its SLO targets. This allows the remaining resources to be distributed among BE jobs, thus improving system utilization and the performance of those services.
By being aware of workload QoS targets, a controller may locate LLC resources in a way that decreases the overall power consumption of a system through the management of performance slack. This is what nejat_cooperative_2022 (nejat_cooperative_2022) propose with their Cooperative Slack Management scheme. It identifies opportunities to generate slack with minimal use of energy. This slack may later be consumed at times of less power-efficient operation and thus increase energy efficiency over time.
Sturgeon (pang_adaptive_2021) is a multi-level resource management system that manages BE and LC workload placement at the cluster level in addition to providing machine-level resource management with a focus on achieving higher performance while not increasing energy consumption. They propose using machine learning models to predict both the performance and power consumption of heuristic changes to resource partitioning based on performance slack. This is to ensure services stay within SLO requirements and that the system does not breach power usage constraints.
Avalon (chen_avalon_2019) is a resource management scheme proposed by chen_avalon_2019 that attempts to improve system utilization by providing “just enough” resources to LC services while ensuring QoS targets are met. This is done by predicting the resource requirements for each incoming LC workload query using an ML model and allocating resources appropriately. The execution time slack of queries is monitored and heuristic adjustments to resource allocation are made if necessary to meet QoS targets. This improves utilization by providing co-located BE jobs with increased resources while ensuring LC QoS.
5.3.5. Summary
Being application-aware means a higher configuration overhead for controllers managing a heterogeneous set of workloads. However, it does allow controllers to optimize the resource partitioning of a system while maintaining workload SLO targets. Publications covered in this section have seen improvements in LC workload throughput by $31.7\%$ (zhao_rhythm_2020), system utilization by $28.9\%$ (chen_avalon_2019), and serverless function execution time by $11\%$ (jeatsa_casy_2022). It is therefore clear that application-aware management of QoS enforcement features may greatly benefit execution environments not already configured to limit congestion over shared resources.
The problem of shared resource management is approached in the context of virtual network functions by (li_rldrm_2020; chintapalli_restrain_2022; tootoonchian_resq_2018; chintapalli_ravin_2023). It is common that such functions are constrained by strict QoS requirements, and are thus traditionally deployed on machines with relatively low total utilization (li_rldrm_2020). By guaranteeing scarce resource access, utilization of such machines may be greatly improved for reduced cost of ownership and increased efficiency.
By observing LC workload deadlines and by efficiently managing performance slack, efficiency (nejat_cooperative_2022; pang_adaptive_2021), server utilization (zhu_dirigent_2016; chen_parties_2019; chen_avalon_2019), or system performance (lo_heracles_2015) may be improved. This is done by identifying times at which slack may be increased at a low cost such that it may be consumed at a later time when the cost is higher.
6. Discussion
Researchers have over the years leveraged the quality of service enforcement features of modern processors to propose a wide range of state-of-the-art proposals that optimize system performance, utilization, fairness, power consumption, and much more. As illustrated in Figure 7, the interest and utilization of such features remains high as more and more creative use cases are found. This is not exclusive to academic circles but includes the industry actors, as indicated by the many collaborative works published between the academic and private sectors. However, the taxonomy presented in this survey shows a distinct lack of publications presenting modern, generally applicable, general-purpose cloud computing approaches. This is especially true when considering papers originating from the public clouds and their providers, who stand to gain major improvements in operational efficiency through such research.
We have found a few exceptions to this trend. Firstly, CacheSlicer (shahrad_provisioning_2021) is a proof-of-concept of an LLC-aware extension of the Microsoft Azure virtual machine scheduler. Secondly, Zeus (zhang_zeus_2021) attempts to improve upon Kubernetes resource management and increase cluster utilization at NetEase. Finally, the intent-driven orchestrator (metsch_intent-driven_2023) tries to learn how changes in resource allocations affect service metrics and then make appropriate adjustments.
This shortage of published material may suggest one of a number of possible scenarios:
  1. (1)
    Larger corporations and cloud providers do not see potential in the available hardware-based QoS platform features and do thus not wish to invest resources into exploring a dead end.
  2. (2)
    The private sector is researching, investing in, and developing efficient ways of utilizing recent QoS features behind closed doors. Thus gaining a competitive advantage by not sharing their findings.
Given the large number of published works surveyed in this paper that show significant improvements to be gained through the effective usage of QoS features, we find scenario (1) highly unlikely. When considering that chip manufacturers have continued to develop and improve, generation by generation, upon a feature that potentially none of their major corporate customers utilize, scenario (2) seems increasingly more likely to reflect reality.
6.1. Privacy Concerns of Private and Public Clouds
The type of cloud operated has a big impact on the type of approaches usable in a given scenario. A private cloud, hosted on-premise, running only software written in-house has vastly different requirements than a cloud running customer software. In the first scenario, cloud operators are able to monitor applications in any way imaginable. This includes prior-to-deployment offline profiling of software. This allows resource management controllers with full application-aware integration to be employed, as leveraged by the intent-driven orchestrator (metsch_intent-driven_2023). However, companies hosting other organizations’ software through infrastructure-as-a-service (IaaS) solutions are required to take additional security measures into account as detailed by shahrad_provisioning_2021 (shahrad_provisioning_2021). In such scenarios, as in the context of public cloud providers, there is no alternative but to treat customer software as black boxes. Nonetheless, managed platform-as-a-service (PaaS) offerings could provide software add-ons allowing customers to provide a resource manager with application-level metrics and targets. Such solutions would make customer software a little more “grey” without becoming an intellectual property protection concern.
Additionally, there may be sensitive data managed by the application that must be handled with discretion due to legal compliance reasons, such as with the EU’s General Data Protection Regulation (GDPR) (noauthor_general_nodate).
6.2. Consolidation of Workloads
Many of the workload consolidation approaches discuss consolidation from the point of view that workloads are either latency-critical (LC) continuously running services or best-effort (BE) jobs. The general idea is to ensure SLOs are met by one (lo_heracles_2015; zhao_rhythm_2020) or more (chen_parties_2019; patel_clite_2020; nikas_dicer_2019; qin_locpart_2017; qiu_slo_2021; li_rldrm_2020; zhu_dirigent_2016; chen_avalon_2019) LC services while utilizing additional machine resources to pack and maximize the performance of BE jobs. This way, the utilization of hardware can be increased while performance expectations of services are met. While the way different types of workloads are used to improve system utilization is sound, the approach assumes a decently static amount of LC traffic and a steady stream of BE jobs awaiting processing. They do not consider scenarios where LC traffic may vary enough that replication scaling is preferred or required, as is common in cloud-based environments. Nor is there any consideration for the underprovisioning that would emerge if no BE jobs are scheduled.
Even though the discussion contexts of these papers are that of cloud computing and macro-scale data centers, the proposed controllers are inherently micro-scale and operate on the scale of a single machine. Therefore, we see a potential for further research, where workload consolidation is tackled in a more holistic fashion at the macro-scale. Where service replication scaling is considered and the notion of LC and BE workloads is approached in a less static way.
6.3. Classification of Workloads
As may be observed by tables 4 and 5, many publications include some form of workload classification as part of their proposed solutions. This is a reasonable way of tackling the resource partitioning problem due to the constraint of finite classes of service (CLOS) available in hardware and subsequently the resctrl resource management interface. Therefore, workloads are required to be generically grouped together according to some scheme. This may range from a simple binary distinction between resource-sensitive and insensitive (e.g. CA (pons_improving_2018)) to more sophisticated program phase classification schemes (e.g. dCat (xu_dcat_2018)).
6.4. QoS Enforcement Feature Strengths and Weaknesses
Given the numerous studies that have successfully applied both cache allocation and memory bandwidth limitation to achieve their respective goals (be it performance, fairness, energy efficiency, utilization, or a combination of the aforementioned goals), it is clear that the hardware QoS enforcement features surveyed in this article are appropriate tools to utilize when optimizing for such goals. Additionally, the aforementioned features have also been shown to be effective in mitigating other issues such as security threats (liu_catalyst_2016).
However, the fundamental indirect nature of QoS enforcement features makes them difficult for practitioners to utilize effectively. In many cases, performance is left on the table that could be unlocked by integrating QoS enforcement feature management into systems such as VM hypervisors and container orchestrators. Such integrations would allow users who are unaware that their hardware supports such tuning to take full advantage of their hardware, automatically and effectively.
7. Research Gap Analysis
As we discussed in Section 6, it is highly likely that public cloud providers make extensive use of the QoS enforcement features found in recent server processors. However, our survey of the academic landscape has made it clear that there is space for holistic solutions that are able to intelligently manage cache and memory bandwidth in the context of multiple machines or even entire data centers. Only a few publications approach QoS management in the context of data centers (shahrad_provisioning_2021; jeatsa_casy_2022), which focus solely on cache partitioning. This leaves room for additional research that could improve the utilization of hardware in smaller, on-premises or edge deployments that generally cannot benefit from the advancements made behind public cloud providers’ closed doors. A fitting approach would be to extend service orchestrators such as Kubernetes (the_kubernetes_authors_kubernetes_2022) with intelligent cache and memory bandwidth awareness.
Observing the evaluations listed in Table 3, it is notable that no evaluation focuses on the specifics of AMD’s QoSE implementation of hardware QoS enforcement features. This is significant as studies have shown inconsistent behavior between hardware generations (sohal_closer_2022), even from the same manufacturer (Intel). Thus, a systematic evaluation where the behavior of AMD’s implementation should be conducted. This would be nicely followed by a comparison of the behaviors of the two chip manufacturers’ implementations.
Additionally, it may be observed that there exists no generic solution to controlling systems spanning different processor models, generations, or even manufacturers. Given the greatly varying degrees of control and feature support between chips, the adoption of such features in general resource management systems may be cumbersome due to the wide range of models to support. Therefore, a general abstraction model may be developed that makes efficient use of the underlying features while providing an interface that is generic and platform-independent. Such an abstraction would lower the friction of feature integration into container orchestrators, VM schedulers, and similar platform-heterogeneous systems.
Furthermore, by assuming that most LC cloud workloads are allocated more resources than required to account for sudden spikes in load without breaking SLOs, an opportunity to improve system-wide efficiency by utilizing this gap between allocation and actual utilization may be observed. Limiting workloads that operate with some margin to their SLOs with intelligent management of shared resources should allow net efficiency gains to be made without impacting the end-user experience. This concept is in part addressed by pang_adaptive_2021 (pang_adaptive_2021). However, their approach only rudimentarily utilizes cache partitioning without touching memory bandwidth limiting.
Finally, tuning multiple enforcement features simultaneously is difficult without experiencing undesired side effects. The indirect nature of QoS enforcement means that the tuning of one workload often affects the performance of co-located workloads. Thus, there exists an opportunity to develop a solution that allows for the tuning of multiple features in unison without negatively affecting neighboring workloads. Such a solution should make use of application awareness to monitor key performance metrics such that the desired outcome is achieved.
8. Concluding Remarks
We have conducted a broad survey of the research landscape and identified how academics are utilizing hardware-based quality-of-service enforcement features implemented in commodity x86 processors. Through our use of a reference snowballing approach, a search space consisting of $3685$ references was reviewed. Out of which, $79$ papers were identified to meet our specified inclusion criteria. We proposed a novel taxonomy of the included papers that structures and classifies the current approaches used by scholars. By observation of this taxonomy, conclusions could be drawn regarding the current general directions of research. Furthermore, gaps in current state-of-the-art research were identified and detailed.
Acknowledgements.
The authors would like to thank Ravi Iyer (Intel Labs) for his input and invaluable feedback. Financial support has been provided in part by the Knut and Alice Wallenberg Foundation under grant KAW 2019.0352 and by the eSSENCE Programme under the Swedish Government’s Strategic Research Initiative.
[…]